# MoveScanner: Analysis of Security Risks of Move Smart Contracts

### v2

@@ 3.3. Privilege Leakage Detection
MoveScanner’s security analysis module includes five essential algorithms. resource leak detection, unchecked return value detection, arithmetic overflow detection, cross-module security detection, and capability leak detection. The resource leak detection, as shown in Listing 2, utilizes control flow and data flow analysis to track the creation, movement, transfer, and destruction of resources across all potential execution paths. It pays special attention to conditional branches and early returns(torlak2010effective, ). Unchecked Return Value Detection checks whether function call return values are properly handled, focusing on Boolean or non-empty returns, shown in Listing 1. Arithmetic Overflow Detection examines addition(schuite2000integer, ), subtraction, multiplication, and ensures that division and modulo operations are handled carefully to avoid potential overflow issues or risks related to division by zero, as shown in Listing 3. Cross-module security detection examines the calls made between different modules, identifying public functions that modify other modules’ global state without proper access control, shown in Listing 4. Capability Leak Detection focuses on the specific capabilities associated with each movement pattern, tracking transfers of capability resources and identifying potential privilege escalation when transferred to untrusted recipients or exposed in public functions, shown in Listing 5.
These five algorithms integrate control flow and data flow analysis, type resolution, and cross-module call analysis, providing comprehensive security verification for Move smart contracts, ensuring safe resource management, return value handling, arithmetic operations, cross-module interactions, and capability transfers.

### v3

@@ 3.3. Privilege Leakage Detection
MoveScanner’s security analysis module includes five essential algorithms. resource leak detection, unchecked return value detection, arithmetic overflow detection, cross-module security detection, and capability leak detection(Wang et al., 2024a). The resource leak detection, as shown in Listing 2, utilizes control flow and data flow analysis to track the creation, movement, transfer, and destruction of resources across all potential execution paths(Mperejekumana et al., 2024). It pays special attention to conditional branches and early returns(Torlak and Chandra, 2010). Unchecked Return Value Detection checks whether function call return values are properly handled, focusing on Boolean or non-empty returns, shown in Listing 1. Arithmetic Overflow Detection examines addition(Schuite et al., 2000), subtraction, multiplication, and ensures that division and modulo operations are handled carefully to avoid potential overflow issues or risks related to division by zero, as shown in Listing 3. Cross-module security detection examines the calls made between different modules, identifying public functions that modify other modules’ global state without proper access control, shown in Listing 4. Capability Leak Detection focuses on the specific capabilities associated with each movement pattern, tracking transfers of capability resources and identifying potential privilege escalation when transferred to untrusted recipients or exposed in public functions, shown in Listing 5.
These five algorithms integrate control flow and data flow analysis, type resolution, and cross-module call analysis, providing comprehensive security verification for Move smart contracts, ensuring safe resource management, return value handling, arithmetic operations, cross-module interactions, and capability transfers(Yin et al., 2024).